OpenStudy (anonymous):

Tim and Tom are trying to earn money to buy a new game system over a 3-month period. Tim saved $45.73 each month. If they need a total of $213.24 to buy the game system, how much does Tom need to earn each of the 3 months in order to buy the game system? A:$137.19 B:$167.51 C:$76.05 B:$25.35

OpenStudy (anonymous):

\[45.73m =213.24\]

OpenStudy (anonymous):

Get m (months) by itself.

OpenStudy (anonymous):

\[m =\frac{ 213.24 }{ 45.73 }\]

OpenStudy (anonymous):

What does m equal?

OpenStudy (anonymous):

one moment

OpenStudy (anonymous):

Oh shoot hang on

OpenStudy (anonymous):

4.66404199475

OpenStudy (anonymous):

Okay nevermind lol. Sorry. \[45.73(3) + t(3) = 213.24\]

OpenStudy (anonymous):

Solve that instead.

OpenStudy (anonymous):

Know what you need to do?

OpenStudy (anonymous):

no! lol

OpenStudy (anonymous):

Get t (toms earned money) by itself

OpenStudy (anonymous):

Whats 45.73 x 3?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

im bad at math

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

45.73 x 3...

OpenStudy (anonymous):

137.19

OpenStudy (anonymous):

213.24-137.19=?

OpenStudy (anonymous):

76.05

OpenStudy (anonymous):

76.05 divided by 3 = ?

OpenStudy (anonymous):

what time is it my computer want tell me 25.35

OpenStudy (anonymous):

t= 25.35. So, Tom would have to earn $25.35 for three months.

OpenStudy (anonymous):

and 8:17 pm for me

OpenStudy (anonymous):

Georgia

OpenStudy (anonymous):

you Georgia

OpenStudy (anonymous):

It's 8:17pm then Anyway, thats your answer